# SAB3R: Semantic-Augmented Backbone in 3D Reconstruction

### v1

@@ 1 Introduction
Current 3D open-vocabulary segmentation methods [51, 42, 68] typically assume access to complete, high-quality point clouds—an assumption that rarely holds in real-world embodied AI scenarios. One major challenge lies in the high cost and complexity of curating large-scale 3D open-vocabulary datasets, even with prior efforts such as ScanRefer [9] and ReferIt3D [1], which remain limited in both scale and diversity. Additionally, existing methods either depend on precise camera poses and sensor calibration for accurate point cloud reconstruction, an impractical requirement in continuously changing environments, or rely on test-time optimization techniques [37, 24], which are computationally expensive and unsuitable for real-time applications. Despite these challenges, human perception effortlessly integrates 2D visual semantics with 3D structural understanding, leveraging depth cues and object motion over a lifetime of interaction [26]. Thus, we aim to explore how a model can simultaneously perform 2D semantic understanding and 3D reconstruction, bridging the gap between segmentation and spatial reasoning in open-vocab settings.
@@
Therefore, our work addresses this challenge by drawing inspiration from human visual perception. As human can seamlessly interpret images by combining 2D visual information with an intuitive understanding of 3D structure. While existing methods take in posed RGB-D sequences or pre-scanned environments, we propose using unposed video as input—a natural and accessible modality for embodied agents operating in real-world settings. As illustrated in Figure 1, our Map and Locate task jointly constructs a 3D geometric map and segments objects specified through open-vocabulary queries. This approach enables simultaneous spatial mapping, semantic understanding, and segmentation without requiring pre-processed point clouds. To this end, we introduce a simple yet effective baseline SAB3R , as shown in Figure 2, which takes unposed images as input and predicts a point map, dense CLIP features, and dense DINOv2 features in a single forward pass.

### v2

@@ 1 Introduction
Current 3D open-vocabulary segmentation methods [51, 42, 68] typically assume access to complete, high-quality point clouds—an assumption that rarely holds in real-world embodied AI scenarios. One major challenge lies in the high cost and complexity of curating large-scale 3D open-vocabulary datasets, even with prior efforts such as ScanRefer [9] and ReferIt3D [1], which remain limited in both scale and diversity. Additionally, existing methods either depend on precise camera poses and sensor calibration for accurate point cloud reconstruction, an impractical requirement in continuously changing environments, or rely on test-time optimization techniques [37, 24], which are computationally expensive and unsuitable for real-time applications. Despite these challenges, human perception effortlessly integrates visual semantics with 3D structural understanding, leveraging depth cues and object motion over a lifetime of interaction [26]. Thus, we aim to explore how a model can simultaneously perform semantic understanding and 3D reconstruction.
@@
To address this challenge, we propose a new approach that takes unposed video as input—a natural and accessible modality for embodied agents operating in real-world settings. Unlike existing methods that rely on posed RGB-D sequences or pre-scanned environments, our method requires neither pre-computed point clouds nor precise camera calibration. As illustrated in Fig. 1, we introduce the Map and Locate task, which jointly constructs a 3D geometric map and segments objects specified through open-vocabulary queries. This formulation enables simultaneous spatial mapping, semantic understanding, and object localization from raw visual input. To this end, we present a simple yet effective baseline, SAB3R , shown in Fig. 2, which takes unposed images as input and predicts a point map, dense CLIP features, and dense DINOv2 features in a single forward pass. We present qualitative results for 3D Reconstruction and Map and Locate task in Fig. 1.